On a still -10°C day, an air film coats the inside of a single glazed glass (k= 0.84 Wm^-1 K ^-1) window. The window is 75 cm by 1m has 3mm thick glass . Assume h= 10. The inside temperature is 20°C . Find the rate of thermal energy flow .


1
Expert's answer
2021-02-02T09:32:00-0500

Rate of flow of thermal energy is, "\frac{dQ}{dt }= \frac{KA\Delta T}{l}"

Where A is the area of the window,

K is the thermal conductivity,

l is the thickness of the window,

"\Delta T" is the temperature difference.


"A = 0.75 \times 1 = 0.75 m^2"

"K = 0.84 Wm^{-1} K ^{-1}"

"\Delta T = 30^\circ C" and "l = 3 mm = 3 \times 10^{-3}m"


Putting all the values, "\frac{dQ}{dt }= \frac{KA\Delta T}{l} = \frac{0.75 \times 0.84 \times 30}{3\times 10^{-3}} = 6.3 KJs^{-1}"